OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (agreene):

start by distributing the 3 and -2 into the parenthesis and you get: 3x+3 = -2x+2+6 combine like terms 3x+3 = -2x+8 add 2x and subtract 3 from each side 3x+2x = 5 combine like terms 5x = 5 divide by 5 x=1

OpenStudy (angela210793):